Given that the risk-free rate is 5%, the expected return on the market portfolio is 20%, and the standard deviation of returns to the market portfolio is 20%, answer the following questions: c. Now suppose that you want to have a portfolio, which pays 25% expected return. What is the weight in the risk free asset and in the market portfolio? d. What do these weights mean: What are you doing with the risk free asset and what are you doing with the market portfolio? e. What is the standard deviation of the portfolio in e? f. What is your conclusion about the effect of leverage on the risk of the portfolio?
Answer #1
Return Std dev
Risk free asset 5% 0%
Market return 20% 20%
Let weight of market return is x and weight in risk free asset is 1-x
20x+5*(1-x) = 25
x = 20/15
x = 1.333
1-x= -0.333
Weight in risk free asset = -0.333
These weights mean that we are borrowing at the risk free rate
And investing in the market portfolio from funds borrowed at risk free rate
Std dev of portfolio 26.67%
The risk of the portfolio would increase based on leverage